OpenStudy (anonymous):

If g(x) = x^2 + 3, find g(4). A: 11 B: 19 C 16 D: 8

OpenStudy (anonymous):

ill fan & medal

OpenStudy (anonymous):

g(x) = (4) ^2 + 3 g(x) = 16 + 3 g(x) = 19 = B

OpenStudy (anonymous):

thank you, i was doing it right just second guessed myself :)

OpenStudy (anonymous):

okayy :)